Ben is three times as old as Carl now. In two years time, Ben will be twice as old as Carl. How old is Carl now?
Answer by jorel1380(3719)

b=3c
b+2=2(c+2)
3c+2=2c+4
c=2
b=6
Carl is 2 now..